# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ 6.1. Literature review over existing guidelines or design recommendations
[…]
| Study | Target audience | Target domain | Design considerations/recommendations/criteria |
[…]
| Guideline | Definition | Purpose | Example |
[…]
| (EASY) G1: Allowing Controlling and Changing Things | How much users can control and change things during adaptation. | A good mix between letting users control things and having the system do some stuff automatically. | |
[…]
| (EASY) G2: Helping Users Easily Use Adaptive Interfaces | Help users to use adaptive interfaces more easily. | Make sure users know how adaptive features work, what they can do, and how to use them properly. | |
| G3: Aligning Adaptations with User’s Chronic Disease | Consider multimorbidity, the understanding, and the progression of their chronic diseases. | Ensure that any adaptations made do not hinder but rather improve users’ ability to effectively manage their chronic condition. | ”Including relevant icons and visuals related to chronic diseases.” (IP 1) ”The process of adaptation needs to consider various chronic diseases, since individuals with certain types of disease may have varied preferences for adaptation.” (From the survey data analysis) |
| (EASY) G3: Considering Users’ Chronic disease | Look at if users have multiple health issues, how well they know their conditions, and how they handle their health problems. | Make sure that any changes do not make it harder but instead help users better manage their health condition. | |
| G4: Aligning Adaptations with App Usage Patterns | Ensure that adaptations align with the frequency, duration, and effort of usage for specific functions. | Improve the effectiveness of implemented adaptations, enhance user experience, and engage users without disrupting their typical app usage. | ”Tasks that require minimal effort and are performed daily may not require manual control from users.” (From memos) ”Users may not desire frequent adaptation to occur if it is a function they will use everyday” (From memos) |
| (EASY) G4: Following the App Use behaviour | Making sure that adaptations match how often, how long, and how much effort users put into using specific functions. | Create a smooth and engaged user experience by having the adaptations. | |
| G5: Aligning Adaptations with User’s Coping Style | Ensure that adaptations are in line with the coping strategies users use to deal with stressful situations. | Adapt the adaptations based on varying health statuses can enhance user engagement with the mHealth application and improve self-management of chronic disease. | ”I prefer frequent and clear reminders for any negative measurement. I do not see the point in avoiding the bad measurement.” (IP 5) ”I haven’t been using the app recently because I have not experienced any negative effects. I prefer not to see my high blood pressure measurement daily because I am afraid of seeing the bad measurement.” (IP 4) |
| (EASY) G5: Matching how users handle stress | Make sure that the changes made to the app match how users handle stress that casued by different health situations. | Tailor the features according to different health conditions can make the mHealth app more engaging and help users better manage their chronic diseases. | |
| G6: Collaborative Adaptation between Users and Caregivers | A cooperative approach in which both primary users and their caregivers work together to customise and fine-tune various aspects of the application, or even jointly utilise the application. | Emphasise the significance of the caregiver’s role throughout the adaptation process and explore strategies for integrating the caregiver’s contributions into the adaptation procedures. | ” I have to look over the health data, understand the app features, and set up some reminders in the app for my family members.” (IP 1,3,4, FP 9) ”I have to set up the app for my parents if the app has provided some flexibility in the functionality or other aspects.” (IP 4) |
| (EASY) G6: Collaboration Between User and Caregiver | Both patients and caregivers work together using the digital app. | Highlight the important role caregivers play in the adaptation process and find ways to include their input in the adaptation procedures. | |
| G7: User Involvement in Adaptations-Assessing User Capability | Assess whether users are physically or mentally capable of handling additional burdens imposed by the system during adaptation. | Avoid overloading users and causing harm through different adjustments. | ”My parents would never consider the option of adapting the app because it is beyond their ability.” (IP 1) ”When I first start using an app, I always customise everything that can be customised. ” (FP 7) |
| (EASY) G7: Considering User’s Ability in Participating in the Adaptation | Check if users can handle additional tasks from the system during adaptation, both physically and mentally. | Ensure that users do not feel overwhelmed or harmed by introducing various changes. | |
| G8: User Involvement in Adaptations-Assessing User’s Willingness | Users exhibit different degrees of engagement in adaptation, which can be shaped by personality, cultural background, and other contextual factors. | Empower users by giving them the choice to either actively participate in shaping the adaptation or passively accept the adaptation. | ”I customise all apps, I will explore how I can adapt the app to my needs (Active involvement).” (FP 7) ”I haven’t even looked into if I can adapt the app or not (Passive involvement).” (FP 2) ”I don’t want to understand or be told what the adaptation is (Passive involvement).” (FP 1) |
| (EASY) G8: Considering User’s Willingness in Participating in the Adaptation | Some users are more involved in making changes to the app than others. This can be influenced by their personality, culture, and other factors. | Give users the option to either actively shape the adaptation or simply accept it as is. |
| Guideline | Definition | Purpose | Example |
| G9: Adjusting Granularity of the Adaptation | Degree or level of detail at which adaptations can be made to an interface. | Strike a balance in granularity to enable users to make meaningful adaptations to the user interface without overwhelming them with extensive adaptations or fatiguing them with minimal adaptations. | ”Once I understand what I am doing, I might want to push that button and get that next level of understanding(adaptation).” (IP 3) ”I think all adaptations should be implemented when users initially log into the app and change the whole app to user needs.” (FP 4,6) |
| (EASY) G9: Changing the Scale of Adaptation in the Interface | The extent to which changes can be made in an interface. | Ensure that users can make meaningful changes to the user interface without overwhelming them with too many changes or leaving them uninterested with too few changes. |
6.2. Evaluation of our Guidelines
We carried out a preliminary evaluation of our guidelines, primarily concentrating on assessing their clarity and perceived usefulness. Our objective was to determine whether target guideline users could effectively comprehend and distinguish between and make use of our guidelines. Given the possibility of limited documentation on design considerations and guidelines, particularly in domains involving new technology, it is imperative to maintain user involvement throughout the process (Hermawati and Lawson, 2015; Zaphiris et al., 2007; Lechner et al., 2013; Zaphiris et al., 2009). Although the guidelines were originally created for developers, designers, and researchers, it is crucial to ensure their organisation, usability, and significance for end users. (Lechner et al., 2013; Zaphiris et al., 2009). Therefore, we strived to further validate the efficacy and acceptance of these guidelines with the same cohort of users involved in our interview and focus group study, aiming to uncover any challenges that may not have surfaced previously and to determine whether we have accurately interpreted their perspectives, presenting them in a way that avoids misunderstandings. We reached out to all participants via email, inviting them to complete a feedback survey. Before administering the feedback survey, participants were briefed on the findings of the user study and provided with an overview of the generated guidelines. To maintain anonymity and streamline the process, participants were required to enter the unique withdrawal code provided during their initial user study registration, allowing us to correlate their responses with both the feedback survey and the user study. Subsequently, participants were requested to assess each of the guidelines and offer suggestions regarding additions, removals, or edits for each guideline. In total, 20 out of 22 participants responded to our feedback survey.
For the purpose of clarification, we designate the original versions of the guidelines as ”GO” (Guideline Original) and the updated versions as ”G”. Figure 11 shows the development of the structure of the guidelines. Initially, the framework table contained seven guidelines organised into three columns: Guidelines, Definition, and Examples. Upon analysis of the qualitative data, we integrated the purpose of the guidelines while reformulating them, leading to a revised structure with Guidelines, Definition, Purpose, and Examples. To enhance clarity, we introduced an ”EASY” version of the guideline alongside the existing guidelines, rather than removing or altering technical terms. Two guidelines (GO2 and GO5) were subdivided into more detailed ones. One guideline (GO6) was removed because it offered limited guidance and overlapped with other guidelines. These modifications resulted in a framework comprising nine guidelines, as detailed in Table 10.
6.3. Clarity of the guidelines
Although most participants found our guidelines understandable (see Figure 12), a few encountered difficulties with certain technical terms, particularly ”Granularity” (GO7) and ”Autonomy” (GO1). These terms may require further clarification or simplification to ensure broader comprehension among users. One participant in particular found it challenging to understand the guidelines, which significantly affected her ability to provide feedback. As a result, she refused to complete the feedback survey and commented:
[…]
@@ 6.3. Clarity of the guidelines
Meanwhile, a critique regarding the difficulty in relating to the examples provided for each guideline underlines the need for clearer and more context-specific illustrations. Participants expressed confusion about how the examples applied to their own experiences and situations, particularly in the case of the Alignment of the adaptation (GO3) guideline (FP 4,7). This discrepancy in understanding could be attributed to the diverse backgrounds and perspectives of the participants. The challenge of creating relevant and accessible examples is a common issue in guideline development, as noted in the literature (Ning et al., 2019). To improve the comprehension of the example, we extracted the sample from our user study as shared by the participants themselves (see Table 10).
Rather than simply replacing terms and expressions with plain language, we acknowledge that achieving clarity and understanding across diverse user groups with varying domain knowledge may pose challenges. Research has explored how different domains can influence individuals’ perspectives on human values in eHealth applications (Obie et al., 2021). It is essential to recognise that the guidelines were crafted primarily for designers experienced in application development, whereas users may perceive and value aspects differently based on their unique perspectives and experiences. Thus, tailoring communication and guidance to align with users’ diverse viewpoints and preferences becomes imperative for effective implementation and user engagement. Therefore, we introduced an additional version of each guideline labelled as ”EASY” in Table 10. This version utilises simple language to describe the name, definition, and purpose of each guideline.
Apart from the clarity of certain guidelines, some guidelines like Alignment of the adaptation (GO2) and User involvement in the adaptation (GO5) have attracted confusion because they can be too broad and difficult to interpret (IP 1,2,6, FP 7,9). The necessity of subdividing these two guidelines into smaller and more granular components is evident. This need is further supported by Figure 12, which illustrates that GO2 and GO5 received the fewest strong agreements on their clarity. Consequently, we have divided the Alignment of the adaptation (GO2) guideline into (see Table 10):

### v2

@@ 6.1. Literature Review over Existing Guidelines or Design Recommendations
A review of general design guidelines for mHealth applications was conducted to establish a foundational understanding of best practices. In 2019, the Xcertia guidelines were introduced to specifically address usability concerns in the mHealth application domain (Board, 2019). The guidelines consist of ten distinct sections that address key usability issues for mHealth applications, as summarized below: 1) Visual design, 2) Readability, 3) App navigation, 4) Onboarding, 5) App feedback, 6) Notifications, 7) Alerts and alarms, 8) Historical data, 9) Ongoing application evaluation, 10) Help resources and troubleshooting. Additionally, Conor et al. (2014) developed nine rubrics to help users evaluate the relevance, quality, functionality, and security of medical applications. Complementing these efforts, Stoyanov et al. (2015) proposed the Mobile App Rating Scale, a framework for evaluating mHealth applications. The scale includes four objective quality dimensions: engagement, functionality, aesthetics, and information quality, as well as a subjective quality dimension. The World Wide Web Consortium (W3C) has established widely recognized accessibility standards, which provide comprehensive guidelines for digital and mobile accessibility (W3C, 2015). WCAG 2.0, released in 2008, introduced foundational accessibility guidelines focused primarily on web accessibility. WCAG 2.1, introduced in 2018, added 17 new criteria to address accessibility challenges in mobile applications, low vision support, and cognitive disabilities. WCAG 2.2 (2023) introduced additional criteria targeting further inclusivity. Despite increasing research on accessibility within the realm of mobile applications (Alshayban et al., 2020; Ross et al., 2017), and some initiatives that address mHealth applications (Milne et al., 2014; Daihua et al., 2015), significant accessibility challenges persist in these applications (Alshayban et al., 2020; Ross et al., 2017). In particular, there is still a lack of comprehensive guidelines to improve the accessibility of mHealth applications. Current initiatives emphasize three critical dimensions in improving accessibility for mHealth applications: 1) Accessible content presentation ensures that content and information are effectively delivered to all users, regardless of their abilities (Radcliffe et al., 2021; Zhou et al., 2020; Stowell et al., 2018). 2) Inclusive interaction focuses on enabling users to seamlessly interact with the application, regardless of physical or cognitive limitations (Radcliffe et al., 2021; Zhou et al., 2020). 3) Reliable and assistive functionality ensures that users can depend on the application to provide accurate and error-free information through assistive technologies (Radcliffe et al., 2021; Zhou et al., 2020).
Chronic diseases remain significantly underrepresented in accessibility research (Mack et al., 2021, 2022). There is a lack of standardized guidelines for tailoring applications related to various chronic diseases for different usage contexts. However, three recurring themes have emerged as critical for the design of AUI in this domain: 1) Transparency emerged as a dominant theme in multiple studies. The researchers consistently underscored the need for transparency in various aspects of AUI, including data management, utilization, and specific decision-making processes (Martin-Hammond et al., 2019; Lage et al., 2014; Mezhoudi et al., 2015). This emphasis on transparency aimed to empower users by providing them with comprehensible explanations regarding the adaptations made by the system. 2) Autonomy is another salient topic that has attracted considerable attention. Numerous studies have highlighted the importance of providing users with a sense of autonomy within AUI, particularly in terms of customizing system content, interaction modalities, and data management procedures (Zhang et al., 2021; Sanders and Martin-Hammond, 2019; Martin-Hammond et al., 2019). This autonomy-centric approach aimed to empower users and cater to their individual preferences and needs. 3) Learnability emerged as a crucial consideration to ensure the accessibility and usability of AUI. Several studies focused on facilitating a smooth and intuitive learning curve for users, especially those with limited technical proficiency. Strategies to improve learnability included the provision of user-friendly interfaces, intuitive navigation systems, and customized information presentations adapted to existing user knowledge and familiarity with relevant concepts (Martin-Hammond et al., 2019; Eslami et al., 2018).
6.2. Evaluation of Guidelines with End-users
A total of 20 participants completed the guideline evaluation survey. Based on their feedback, the original set of guidelines (Version One) was refined to produce a more comprehensive and actionable set (Version Two). As illustrated in Figure 7, two guidelines (G3v1 and G5v1) were expanded into more specific subguidelines to improve clarity and usability. One guideline (G6v1) was removed due to its limited contribution and conceptual overlap with others. These revisions led to a final structure consisting of nine distinct guidelines in Version Two, as shown in Figure 7.
6.2.1. Clarity of the guidelines
@@
Although most of the participants found the guideline understandable, a few encountered difficulties with certain technical terms, particularly ”Granularity” (G7v1) and ”Autonomy” (G1v1). These terms may require further clarification or simplification to ensure greater comprehension among users. Meanwhile, a critique regarding the difficulty in relating to the examples provided for each guideline underlines the need for clearer and more context-specific illustrations. Participants expressed confusion about how the examples were applied to their own experiences and situations, particularly in the case of the alignment of the adaptation (G3v1) guideline. The challenge of creating relevant and accessible examples is a common issue in guideline development, as noted in the literature (Ning et al., 2019). It is essential to recognize that the guidelines were developed primarily for software practitioners, whereas end-users may interpret and prioritize aspects of these guidelines differently due to their distinct experiences, needs, and contextual understanding.
Some guidelines like G3v1 (Alignment of the Adaptation) and G5v1 (User Involvement in the Adaptation) have attracted confusion because they can be too broad and difficult to interpret. This was further supported by the clarity ratings, where both G3v1 and G5v1 received the lowest levels of strong agreement on clarity, each at only 40%. To address this issue, G3v1 was subdivided into more focused and granular guidelines: G3v2 (Aligning Adaptations with User’s Chronic Disease), G4v2 (Aligning Adaptations with App Usage Patterns) and G5v2 (Aligning Adaptations with User’s Coping Style). G5v1 is broken down into: G7v2 (User Involvement in Adaptations-Assessing User Capability) and G8v2 (User Involvement in Adaptations-Assessing User’s Willingness) (see Figure 7). Upon further examination, G6v1 (Timing of the adaptation) has been removed from our guidelines because participants perceive it as too general, affecting its clarity, and it overlaps with G4v1.
6.2.2. Usefulness of the guidelines
When evaluating the guidelines, participants expressed difficulty in relating to some of them and struggled to understand their purpose. It is frequently observed in the development of guidelines that they tend to state actions without explaining their rationale or offering implementation advice (Carter, 1999). We have supplemented each guideline with its purpose to offer context and aid in understanding. Furthermore, participants with experience in application design noted the potential overlap between the proposed guidelines and existing general guidelines for designing mHealth applications. This overlap in guidelines aligns with similar findings from other studies (Srivastava et al., 2021). In addition, the participants in the evaluation study emphasized the need to prioritize certain guidelines under different situations. This emphasizes the challenge of designing for the hypothetical ”general” user (Carter, 1999). Feedback from the survey revealed that users engage with chronic disease-related applications in a variety of ways. For instance, younger users often utilize these applications to independently monitor health metrics, which can render guidelines focused on user-caregiver collaboration, such as G4v1, less relevant for this subgroup. The achievement of all design goals for a computer-based product or service often involves trade-offs (Johnson, 2020), as the guidelines proposed in this study, discussed in Section 4.2, are not exempt from such tensions, with certain recommendations potentially conflicting depending on user needs and system priorities.
6.3. Evaluation of Guidelines with Software Practitioners
A total of 43 software practitioners completed our evaluation survey. Each participant was assigned a unique identifier (e.g., [S1]) to reference their input throughout the analysis. Based on their feedback, guidelines have been refined to better correspond to the mHealth application design, with actionable tips added to help developers apply the guidelines in diverse design contexts. One new guideline is introduced (G9v3), while another is removed due to redundancy and limited applicability (G5v2). The resulting nine guidelines (Version Three) are illustrated in Figure 7, with detailed descriptions for each guideline provided in Section 6.4.
| Demographics # % of Participants Gender Female 13 30% Male 29 70% Age 18 - 24 18 42% 25 - 34 17 40% 35 - 44 5 12% 45 - 54 2 5% 55 - 64 1 2% Company size Less than 10 11 21% 11-50 4 8% 51-100 9 17% 101-500 4 8% 501-1000 5 9% More than 1000 8 15% Prefer not to say 2 4% Years of working experience 0-2 years 17 40% 3-5 years 11 26% 6-10 years 11 26% 11+ years 4 9% Ethnicity simplified White 20 47% Asian 14 33% Black 4 9% Other 3 7% Mixed 2 5% | Demographics # % of Participants Country of residence Australia 11 26% Canada 7 16% Portugal 6 14% Poland 4 9% Mexico 3 7% Chile 2 5% India 2 5% Italy 2 5% United States of America 2 5% Greece, Netherlands, New Zealand, United Kingdom of Great Britain and Northern Ireland 1% Each Roles in the team * Programmer 21 49% User interface or Graphical user interface designer 16 37% Software architect 15 35% Tester 10 23% Project manager 8 19% App animator or operations developer/engineer 4 9% QA engineer 3 7% Requirements analyst 2 5% Business consultant/Marketing manager/Sales personnel, Technical Lead, Researcher 2% Each Experience in health-related applications Chronic disease management (e.g., diabetes tracking apps) 14 33% General health and wellness Tools (e.g., diet apps) 14 33% Specialized applications (e.g., tumor detection) 12 28% Experience in related fields (e.g., testing for health apps) 3 7% |
| * This does not added up to 100%, because some participants took several roles. Other categories of demographic data may not sum to 100% due to rounding. | |
6.3.1. Demographic of software practitioners
Table 5 provides a summary of the demographic details of the participants. Participants are predominantly men (70%) and relatively young, with 42% aged 18-24 and 40% aged 25-34, indicating a sample largely composed of early career professionals. Most of the respondents had 0-2 years (40%) or 3-10 years (52%) of experience, and only 9% had more than 11 years of experience. The largest groups by geography hail from Australia (26%), Canada (16%), and Portugal (14%), with smaller groups from various countries in Europe, North America, and Asia. The participants are employed in organizations of varying sizes, with 21% working in small companies (with fewer than 10 employees) and others in medium to large organizations. In terms of professional roles, the largest group is identified as programmers (49%), followed by UI designers (37%), software architects (35%), and testers (23%). These demographics highlight the diversity of professional backgrounds and responsibilities of survey participants within the software development sector.
The responses reveal a diverse range of experiences in developing health-related applications, highlighting significant contributions to chronic disease management (33%), general health tools (33%), specialized applications (28%) and other experiences in health-related fields (7%). Many survey respondents reported working on applications for managing chronic diseases, incorporating features such as tracking vital signs, medication adherence, and personalized health recommendations. Others contributed to general health and wellness applications, including applications promoting healthy routines, dietary management, and senior-focused medication management systems. Specialized projects ranged from tumor detection applications and mental health solutions to orthodontics-related tools and the digitization of healthcare facilities like nursing homes and hospitals. Participants also played key roles in testing, UI design, and information technology (IT) support.
6.3.2. Understandability and usefulness of guidelines
Figure 8 provides insight into participants’ perceptions of the guidelines related to user adaptation in applications, evaluating both their understanding and their usefulness. Most of the participants reported strong agreement with the clarity of G1v2 (49%) and G2v2 (53%), indicating that these two guidelines were among the most easily understood and distinguishable. Similarly, G3v2 and G4v2 received strong agreement from 51% of the respondents, indicating that they were generally well understood. In contrast, G9v2 and G5v2 had lower strong agreement ratings (37% and 33%, respectively), suggesting that these guidelines were perceived as more ambiguous and may require further clarification. In terms of perceived usefulness, G3v2 stood out as the guideline most positively rated, with 70% of the participants strongly agreeing on its utility, followed by G2v2 (65%) and G1v2 (60%). Guidelines such as G6v2 and G9v2 also received considerable support, with 51% and 44% in strong agreement, respectively. The survey participants overwhelmingly demonstrated a strong preference (91%) for using the proposed guidelines over existing standalone guidelines to design mHealth applications, although a small minority (9%) expressed skepticism. In general, while most guidelines were perceived as useful and understandable, such as G9v2 and G5v2, may require refinement to improve clarity and effectiveness. The positive reception of user-centered guidelines highlights their value in improving usability in applications.
6.3.3. Strengths of the guidelines
Software practitioners also provided several comments that highlighted additional strengths and limitations of the proposed guidelines (see Figure 9). The guidelines demonstrate significant strengths in creating user-centric, personalized applications, particularly for individuals who manage chronic diseases. Personalization emerged as the theme most frequently endorsed, referenced in 78% of participant responses, with many commending the guidelines for enabling a personalized application experience. In particular, G3v2 (Aligning Adaptations with Chronic Diseases) and G1v2 (Promoting User Controllability and Autonomy) were specifically praised for empowering users to manage their health while maintaining the flexibility and user-friendliness of the system. Another key strength is the emphasis on empowerment and autonomy, identified in 62% of the responses. The participants appreciated that the guidelines allow users to control application adaptation, reduce frustration, and encourage sustained use. G1v2 and G9v2 were frequently cited as ensuring user control without overwhelming them. Caregiver collaboration, addressed in G6v2, was mentioned in 38% of the responses, emphasizing its critical role in shared health management. Including caregivers not only supports users who need additional help, but also improves adherence to health routines through clear communication and shared decision making. A focus on context-aware adaptations was observed in 85% of the responses, highlighting the importance of aligning the application features with usage patterns, coping styles, and real-time user behavior and chronic disease management. Guidelines such as G5v2 (Aligning Adaptations with Coping Styles) and G4v2 (Aligning Adaptations with App Usage Patterns) are considered critical to maintaining the usability and relevance of the application.
6.3.4. Limitations of the guidelines
While the proposed guidelines demonstrate substantial strengths, the evaluation also uncovered several limitations that may hinder their practical implementation. One key challenge lies in the difficulty of concretely linking certain guideline elements to specific mHealth application design decisions, with some components perceived as inconsistently applicable. For example, [S4] highlighted the challenge of clearly defining the concept of user involvement and its applicability across various types of mHealth applications. One common concern is the risk of overwhelming complexity, as participants emphasized the challenge of maintaining an appropriate balance between the level of granularity, user autonomy, and overall interface simplicity. [S15] noted:
Guidelines such as G1v2 (Promoting User Controllability and Autonomy in Adaptation) and G9v2 (Adjusting the Granularity of the Adaptation) exemplify this tension, as they must navigate the fine line between offering users sufficient control and customization without introducing excessive complexity that could hinder usability or reduce engagement. An additional limitation identified was the challenge in accommodating the variation associated with chronic disease requirements. Guidelines such as G3v2 (Aligning Adaptations with User’s Chronic Disease) must accommodate fluctuating health condition, which may require frequent updates, risking user fatigue or misaligned adaptations. [S23] noted that the willingness of users to modify the application can vary depending on the context. For example, users might initially resist making changes but become more open to adaptation as their familiarity with the applications or their health condition evolves over time. Privacy and security concerns are also prominent, especially when sharing sensitive health data with caregivers. [S16] highlighted the constant risk of fraud and the possibility of unauthorized access to patient personal medical information. Scalability and long-term usability challenges emphasize the importance of continuous testing and refinement to ensure that the system remains relevant and effective over time. Furthermore, [S6] pointed out a potential drawback of customization.
[S18] noted that if users are not in a suitable physical or mental state, prolonged use of the application can decrease, reducing its effectiveness. Furthermore, [S9] raised concerns about excessive dependency on the application, especially during urgent or emergency situations, suggesting that such features should be limited and paired with clearer usage instructions to prevent misuse. Several participants emphasized that certain guidelines warrant particular attention, with G5v2 (Aligning Adaptations with User’s Coping Style) being especially challenging. Developing an application that effectively reflects the user’s coping style necessitates continuous fine-grained data collection over time, which raises concerns about privacy, security, and the potential for user fatigue from repeated prompts or data input. [S31] illustrated an intriguing situation in which the importance of different guidelines can fluctuate, highlighting the interconnected nature of the guidelines. This participant also noted that specific guidelines can interact with others, implying that changes to one may require adjustments to another.
Response to the limitations. Some limitations noted earlier have been thoroughly examined in the AUI literature, including concerns such as privacy, learnability, high complexity, and usability challenges (Paymans et al., 2004; Gajos et al., 2008; Wesson et al., 2010; Lavie and Meyer, 2010; Peissner and Edlin-White, 2013; Wang et al., 2024). Our aim is to highlight these issues in the guidelines while offering strategies for designing AUIs that effectively address or offset these drawbacks to meet users’ needs. Our attention will be directed towards improving the clarification of the guidelines, connecting this with the design of the mHealth application, and refining the guidelines associated with the management of chronic diseases.
6.3.5. Suggestions for the guidelines
Participants also offered valuable suggestions to improve the practicality and impact of the guidelines. Several respondents emphasized the importance of translating the guidelines into clear, actionable steps for development teams. Guidelines such as G1v2, G3v2, G5v2, G6v2, and G9v2 were identified as needing more concrete implementation strategies, particularly in areas such as adaptive learning pathways and data privacy protections. For example, [S1] suggested improving G1v2 (Promoting User Controllability and Autonomy in Adaptation) by allowing users to gradually learn about controls through tutorials or default settings. Furthermore, [S28] emphasized the incorporation of explicit consent mechanisms for caregiver access under G6v2 (Collaborative Adaptation between Users and Caregivers), as well as the provision of secure sharing options, such as temporary access tokens or granular data permissions for collaborative features. The participants called for clarity on the context-specific guidance for implementation, with a strong focus on understanding the interaction between different guidelines. For example, [S11] highlighted the importance of examining how guidelines interact in scenarios that involve collaborative adaptations with caregivers. Further recommendations included developing a holistic onboarding process to assess the physical and mental capacities of users. This process should allow periodic reassessments to ensure that users are not permanently classified based on their initial evaluations. Participants emphasized the need for more explicit and context-specific guidance to support the effective implementation of the guidelines. [S11] highlighted the importance of examining how guidelines interact in scenarios involving collaborative adaptations with caregivers. Several participants stressed the importance of conducting a continuous and iterative evaluation of the guidelines throughout the actual development process. Such practical evaluations would help identify gaps between the theoretical guidance and its usability in real development environments. In addition, several participants advised to perform extensive user research covering various demographics, such as age, socioeconomic status, and cultural background, to ensure that the guidelines are applicable and inclusive across the board. Some participants also advised to consider offering toolkits or design templates to developers, particularly for complex aspects such as G3v2 (Aligning Adaptations with the User’s Chronic Disease) and G9v2 (Adjusting Granularity of the Adaptation). These could serve as an initial guide for developers who might lack experience in dealing with AUIs. In addition, it could be advantageous to include examples or suggestions for the implementation of the guidelines in multi-user scenarios, such as family-oriented applications or shared interfaces between caregivers and patients.
Response to the suggestions. The guidelines have been refined in response to the specificity of the suggestions in the context of mHealth application design. A new guideline, G9v3 (Ensuring Privacy-Conscious Adaptive Mechanisms), has been introduced to improve practical applicability, while G5v2 (Aligning Adaptations with User’s Coping Style) has been removed. To assist software practitioners, actionable tips have been provided to apply the guidelines in various scenarios. Although some suggestions hold significant promise, they could be further explored and refined in future work.
6.4. Guidelines
| ID | Guidelines | End user | Software practitioner |
| G1v2 | Promoting User Controllability and Autonomy in Adaptation | Important | Important |
| G2v2 | Providing User Support for Adaptation | Critical | Critical |